OpenStudy (anonymous):

find the slope of a line that passes through (-2, -3) and (1, 1) 1/3 1 2 4/3

OpenStudy (anonymous):

Use the slope formula\[m = \frac{ y_2 - y_1 }{ x_2-x_1 }\]where x_2 and y_2 represents one coordinate and the x_1 and y_1 represents the 2nd coordinate. Slope = rise over run right?

OpenStudy (anonymous):

yes its rise over run

OpenStudy (anonymous):

since your point (1,1) is higher than your point (-2,-3) on a graph. use the (1,1) as your x_2, and y_2 and the other point (-2,-3) as your x_1 and y_1

OpenStudy (anonymous):

sooo 

OpenStudy (anonymous):

I don't understand what I do next

OpenStudy (anonymous):

@agent0smith

jimthompson5910 (jim_thompson5910):

\[\large m = \frac{ y_2 - y_1 }{ x_2-x_1 }\] \[\large m = \frac{ 1 - (-3) }{ 1 - (-2) }\] \[\large m = ???\]

OpenStudy (anonymous):

so I do -3 x -3 first I guess which = 9 then 1 - 9 = -8 -2 x -2 = 4 1 - 4 = -3 righttttttttttttttttt????

OpenStudy (anonymous):

then 8 divided by 3?

OpenStudy (agent0smith):

See jim's post, follow it slowly and carefully. Not sure why you're doing -3*-3 and -2*-2, there's no multiplication involved.

OpenStudy (anonymous):

the (-3) in it doesn't that mean your supposed to multiply

OpenStudy (agent0smith):

1-(-3) means 1 minus negative 3

jimthompson5910 (jim_thompson5910):

also, 1 - (-3) is really the same as 1 + 3

OpenStudy (agent0smith):

What makes you think (-3) means you multiply...? Make sure you realize what mistake you are making there...

OpenStudy (anonymous):

ok so 1 - -3 = 4

OpenStudy (agent0smith):

Correct

OpenStudy (anonymous):

1 - -2 = 3

OpenStudy (anonymous):

the reason I thought it was multiplication was because commonly with pemdas and sorts the parenthesis can mean you multiply the things inside

OpenStudy (anonymous):

so 4/3

OpenStudy (agent0smith):

But still... what would you multiply it by (-3).. there is nothing other than a -3 in there. There's nothing to multiply it by.

OpenStudy (agent0smith):

(-3) does not mean (-3)^2

OpenStudy (anonymous):

I thought it was sorta like -3^ with the exponents and I though oh it must multiply by itself

OpenStudy (agent0smith):

Make sure you don't continue thinking that... you will make a lot of mistakes if you do. If there's no exponent, don't assume there's one supposed to be there.

OpenStudy (anonymous):

ya sorry math confuses me haha :) thank you for helping me though I really appreciate it

OpenStudy (agent0smith):

And yes, 4/3 is correct :)